# CHESS : Optimizing LLM Inference via Channel-Wise Thresholding and Selective Sparsification

### v1

@@ 3.3 Efficient Sparse Kernels
(a) Attention projection (spvmm kernel)
(b) Down projection (spvmm kernel)
(c) Up projection (vmmsp kernel)
Figure 5: Average downstream performance and end-to-end speedups of each method under different sparsity levels.
@@
4 Experiments
In this section, this paper first introduces the dataset, comparisons, and implementation details. Then, this paper presents the main results over 8 downstream tasks in terms of the model performance and model efficiency. Besides, this paper also conducts an ablation study across different sparsification module and analysis on efficiency over different sparsity level.
4.1 Datasets and Experimental Setup
Datasets We utilize ARC Challenge (Arc-C), ARC Easy (Arc-E), BoolQ, HellaSwag (HS), OpenbookQA (QA), PIQA, SCI-Q, Winogrande (WG) as benchmarks for downstream tasks, employing the Evaluation Harness library from Eleuther AI to ensure consistency with and Lee et al. (2024). These tasks are designed to assess various aspects of the language model’s performance, including comprehension, common sense, and reasoning abilities, which effectively illustrate the model’s capability loss with activation sparsification.
Comparisons To validate the effectiveness of the proposed CHESS , we conducted experiments using several state-of-the-art LLMs, including Llama-2-7B, Llama-2-13B, Llama-2-70B, Llama-3-8B and Mistral-7B. These models feature different attention mechanisms, specifically MHA and GQA, and utilize SwiGLU as the activation function for the FFN modules. For our main evaluation, we tested four different configurations across all five LLMs:
  • Base Model: the LLM model without any activation sparsification.
  • CHESS w/: the proposed method including channel-wise thresholding and selective sparsification.
For the ablation study, we evaluate the following three models:
  • Llama-3: the Llama-3 8B model without activation sparsification.
Implementation Details For all models involving activation sparsification, thresholds are sampled from a subset of the C4 dataset (Raffel et al., 2020). Following the settings in CATS (Lee et al., 2024), the sparsity level $k$ is set to 0.5, where the accuracy drop is minimal while the inference latency significantly decreases. The proposed method was implemented using the PyTorch v2.2.2 and HuggingFace Transformers v4.39.3. End-to-end decoding speedups are measured on a randomly collected subset of C4 dataset. Kernel efficiency and end-to-end speedup experiments are conducted with FP32 precision on a personal computer equipped with an Intel Core I9-12900K CPU and 64GB of DDR4 memory. Since our work can be applied to quantized models as well, changing weight precision to FP16 or even lower bit-width quantizations does not materially affect our results (Lee et al., 2024).
@@ 4.2 Main Results on Downstream Tasks
Table 1 compares the accuracy of different models across 8 downstream tasks and Figure 3 evaluates the end-to-end inference speedups. Experimental results draw the following conclusions.
Channel-wise thresholding can reduce accuracy degradation while achieving comparable sparsity. The proposed CHESS w/o exhibits a smaller average performance drop of 1.07 across 5 base models and 8 downstream tasks, compared to the 1.70 degradation observed with CATS. Specifically, CHESS w/o consistently outperforms CATS on ARC Easy, BoolQ, and HellaSwag, while showing modest gains on the remaining benchmarks. Moreover, CHESS w/o achieves a comparable sparsity to CATS.
Selective sparsification of attention modules further improves sparsity while maintaining model accuracy. Compared to CHESS w/o, the average performance of CHESS w/ degrades by 0.04 on Llama-2-7B and 0.61 on Llama-3-8B, respectively. Interestingly, for larger models such as Llama-2-13B, Llama-2-70B, and Mistral-7B, CHESS w/ demonstrates comparable or even slightly superior overall performances. More specifically, CHESS w/ outperforms on OpenbookQA, but underperforms on ARC Easy, HellaSwag and BoolQ, while showing similar results on ARC Challenge, PIQA, SciQ, and Winogrande. These results demonstrate that the additional selective sparsification on attention modules has minimal impact on performance. In comparison to CATS, CHESS w/ consistently delivers superior average performance with fewer activated parameters.
@@ 4.3 End-to-End Decoding Speedup
CHESS achieves end-to-end speedups of up to 1.27x compared to Transformers baselines. The proposed CHESS w/ achieves the highest speedup of 1.25x on Llama-2-7B and Llama-2-13B, and 1.27x on Llama-3-8B and Mistral-7B, respectively. When not employing attention sparsification, CHESS w/o achieves comparable speedups to CATS, which is 1.17x on Llama-2-7B and Llama-2-13B, 1.20x on Llama-3-8B, and 1.21x on Mistral-7B, respectively. This is because of the comparable parameters activated per decoding pass of these two methods. Due to the limited capacity of main memory of edge devices, we did not perform the end-to-end speedup test for the Llama-2-70B model. However, based on the activated parameter count per inference pass, its speedup is estimated to be similar to that of Mistral-7B.
@@ 4.5 Kernel Efficiency
As illustrated in Figure 4, this paper conducts a comparative analysis of the latency against sparsity level between the proposed custom sparse kernel and the dense kernel in PyTorch (Paszke et al., 2019). At a sparsity level of 0, the vmmsp kernel used for up projections demonstrates slightly lower latency compared to the PyTorch dense kernel. Conversely, the spvmm kernel, utilized by attention projections and down projections, exhibits slightly higher latencies than the dense kernel. This increased latency is primarily due to the advanced loop tiling and reordering strategies, which cause slight performance degradation at low sparsity levels.
@@ 4.6 Impact on Different Sparsity Levels
Regarding end-to-end speedup, CHESS w/ models exhibit the highest speedup at all sparsity levels above 0.3, attributed to the selective sparsification of attention modules. Specifically, CHESS w/ achieves speedups of 1.46x and 1.72x at sparsity levels of 0.7 and 0.9, respectively, compared to 1.33x and 1.52x for CATS. However, at a sparsity level of 0.3, the CHESS w/ model exhibits slightly reduced speedup, mainly due to the suboptimal efficiency of the spvmm kernel kernel at lower sparsity levels.
@@ 4.7 Extended Comparisons with State-of-the-Art Training-Free Pruning Methods
As presented in Table 3, the proposed CHESS method achieves superior performance in most benchmarks while activating comparable or fewer parameters compared to both Relufication and Wanda. Specifically, CHESS with a sparsity level of 0.7 outperforms other methods on several benchmarks including Arc Challenge, Arc Easy, HellaSwag, OpenbookQA, PIQA, SciQ, and Winogrande. Despite using only 54.92% of the model’s parameters per decoding pass, CHESS delivers an average performance (60.21) that surpasses Wanda (56.41) and Relufication (28.94). These results emphasize the advantage of CHESS over existing methods.
@@ 5 Related Work
Non-autoregressive decoding approaches, such as speculative decoding (Leviathan et al., 2023; Zhou et al., 2023) or Medusa (Cai et al., 2024), seek to convert autoregressive decoding process of LLMs into parallel decoding to mitigate memory access overhead. However, these methods simultaneously impose increased computational demands, which presents significant challenges for deployment on personal devices with limited processing capabilities.
@@ 6 Conclusion
7 Limitations
The limitations of this work can be summarized in two main aspects. First, while CHESS achieves lower accuracy degradation compared to existing methods with fewer activated parameters, it still experiences a noticeable accuracy loss, particularly at higher sparsity levels. Future research could explore fine-tuning techniques to help mitigate this decline in performance.
Second, our method performs optimally with a batch size of 1, which is suitable for edge deployment scenarios where typically only a single user is involved. However, under larger batch sizes, the structured sparsity of activation tensors deteriorates into unstructured sparsity, thereby limiting the potential for end-to-end speedups. This limitation constrains its effectiveness in data center deployments, where larger batch sizes are common.

### v2

@@ 4.1 Datasets and Experimental Setup
(a) Attention projection (spvmm kernel)
(b) Down projection (spvmm kernel)
(c) Up projection (vmmsp kernel)
Figure 5: Downstream performance and end-to-end speedups of each method under different sparsity levels.
@@ 4.2 Main Results on Downstream Tasks
Table 1 compares the accuracy of different models across eight downstream tasks and Figure 3 evaluates the end-to-end inference speedups. Experimental results draw the following conclusions.
Channel-wise thresholding can reduce accuracy degradation while achieving comparable sparsity. Achieving a comparable sparsity, the proposed CHESS w/o exhibits a smaller average performance drop of 1.07 across five base models and eight downstream tasks, compared to the 1.70 degradation of CATS. Specifically, CHESS w/o consistently outperforms CATS on ARC Easy, BoolQ, and HellaSwag, while showing modest gains on the remaining benchmarks.
Selective sparsification of attention modules further improves sparsity while maintaining model accuracy. Compared to CHESS w/o, the average performance of CHESS w/ degrades by 0.04 on Llama-2-7B and 0.61 on Llama-3-8B, respectively. Interestingly, for larger models such as Llama-2-13B, Llama-2-70B, and Mistral-7B, CHESS w/ demonstrates comparable or even slightly superior overall performances. Specifically, CHESS w/ outperforms on OpenbookQA, but underperforms on ARC Easy, HellaSwag and BoolQ, while showing similar results on ARC Challenge, PIQA, SciQ, and Winogrande. These results demonstrate that the additional selective sparsification on attention modules has minimal impact on performance. In comparison to CATS, CHESS w/ consistently delivers superior average performance with fewer activated parameters.
@@ 4.3 End-to-End Decoding Speedup
CHESS achieves end-to-end speedups of up to 1.27x compared to Transformers baselines. When not employing attention sparsification, CHESS w/o achieves comparable speedups to CATS, which is 1.17x on Llama-2-7B and Llama-2-13B, 1.20x on Llama-3-8B, and 1.21x on Mistral-7B, respectively. This is because of the comparable parameters activated per decoding pass of these two methods. When employing attention sparsification, the proposed CHESS w/ achieves the highest speedup of 1.25x on Llama-2-7B and Llama-2-13B, and 1.27x on Llama-3-8B and Mistral-7B, respectively. Due to the limited capacity of main memory of edge devices, we did not perform the end-to-end speedup experiment for the Llama-2-70B model. However, based on the activated parameter count per inference pass, its speedup is estimated to be similar to that of Mistral-7B and Llama-3-8B.
@@ 4.5 Kernel Efficiency
As illustrated in Figure 4, this paper compares the latency against sparsity level between the proposed custom sparse kernel and the dense kernel in PyTorch (Paszke et al., 2019). At a sparsity level of 0, the vmmsp kernel used for up projections demonstrates slightly lower latency compared to the PyTorch dense kernel. Conversely, the spvmm kernel, utilized by attention projections and down projections, exhibits slightly higher latencies than the dense kernel. This increased latency is primarily due to the advanced loop tiling and reordering strategies, which cause slight performance degradation at low sparsity levels.
@@ 4.6 Impact on Different Sparsity Levels
Regarding end-to-end speedup, CHESS w/ exhibits the highest speedup at all sparsity levels above 0.3, attributed to the selective sparsification of attention modules. Specifically, CHESS w/ achieves speedups of 1.46x and 1.72x at sparsity levels of 0.7 and 0.9, respectively, compared to 1.33x and 1.52x for CATS. However, at a sparsity level of 0.3, the CHESS w/ exhibits slightly reduced speedup, mainly due to the suboptimal efficiency of the spvmm kernel at lower sparsity levels.
@@ 4.7 Extended Comparisons with State-of-the-Art Training-Free Pruning Methods
As presented in Table 3, the proposed CHESS method achieves superior performance in most benchmarks while activating comparable or fewer parameters compared to both Relufication and Wanda. Specifically, CHESS with a sparsity level of 0.7 outperforms other methods on several benchmarks including Arc Challenge, Arc Easy, HellaSwag, OpenbookQA, PIQA and Winogrande. Despite using only 54.92% of the model’s parameters per decoding pass, CHESS delivers an average performance (60.21) that surpasses Wanda (56.41) and Relufication (28.94). These results emphasize the advantage of CHESS over existing methods.
@@ 5 Related Work
Non-autoregressive decoding approaches, such as speculative decoding (Leviathan et al., 2023; Zhou et al., 2023) or Medusa (Cai et al., 2024), seek to convert autoregressive decoding process of LLMs into parallel decoding to mitigate memory access overhead. However, these methods impose increased computational demands, which presents significant challenges for deployment on personal devices with limited processing capabilities.
@@ 6 Conclusion
Limitations
The limitations of this work can be summarized in two main aspects. First, while CHESS achieves lower accuracy degradation compared to existing methods with fewer activated parameters, it still experiences a noticeable accuracy loss at higher sparsity levels. Future research could explore fine-tuning techniques to mitigate this decline in performance. Second, CHESS is optimized for inference with a batch size of one, which is suitable for edge deployment scenarios typically involving a single user. However, under larger batch sizes, the structured sparsity of activation tensors deteriorates into unstructured sparsity, limiting potential speedups and reducing effectiveness in data center deployments, where larger batch sizes are common.
Acknowledgements
We thank all the reviewers for their insightful comments. This work is supported by the National Natural Science Foundation of China (No. 62472330, U20A20177, 62272348, U22B2022), the National Key Research and Development Program of China (No. 2022YFB3104502), the State Key Laboratory of Computer Architecture (ICT, CAS) under Grant No. CARCH A202112, and Wuhan Science and Technology Joint Project for Building a Strong Transportation Country (No.2023-2-7).